The sum of three consecutive multiples of 8 is 888. Find the multiples. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the three consecutive multiples of 8 be 8x, 8(x+1) and 8(x+2). According to the question,
8x + 8(x+1) + 8(x+2) = 888
⇒ 8 (x + x+1 + x+2) = 888(Taking 8 as common)
⇒8(3x+3)=888
⇒3x+3=888/8
⇒3x+3=111
⇒3x=111-3
⇒3x=108
⇒x=108/3
⇒x=36
Thus, the three consecutive multiples of 8 are: 8x=8×36=288 8(x+1)=8×(36+1)=8×37=296
8(x+2)=8×(36+2)=8×38=304

Thanks